OpenStudy (anonymous):

intergral of root (1 + (e powered by x)) with respect to x

OpenStudy (experimentx):

\[\int\limits_{}^{}\sqrt{1+e^x}dx\]

OpenStudy (anonymous):

the contents are enclosed in a square root

OpenStudy (anonymous):

that's it, how do i get to the answer ?

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sqrt{1+e^x}}{e^x} \cdot e^x dx \text{ let } e^x=\tan^2(u)\] Haven't tried this yet but it might work

OpenStudy (experimentx):

i am looking for it too.

myininaya (myininaya):

I think will work actually After the sub and writing \[2\int\limits_{}^{}\sec^2(u) \csc(u) du\] you will need to do integration by parts

myininaya (myininaya):

integration by parts just once

myininaya (myininaya):

Let me know if you run into anymore trouble :)

OpenStudy (anonymous):

there was a problem in network, thanks myninaya

OpenStudy (experimentx):

did you solve that?? if you have done it and i would be happy to know

OpenStudy (anonymous):

am still crackin my head to see where this leads me

OpenStudy (experimentx):

you can get the answer from wolf ...but it does not show steps

myininaya (myininaya):

Well I'm still here if you can't follow be recipe :)

myininaya (myininaya):

my*